OpenStudy (anonymous):

Solve Bc 0.1=(Bc)/(100+Bc) Please tell me process

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@zepdrix please help!

zepdrix (zepdrix):

Solve for Bc?

zepdrix (zepdrix):

\[\Large\rm 0.1=\frac{B_c}{100+B_c}\]

zepdrix (zepdrix):

Let's multiply both sides by 10, so we don't have to deal with a decimal value.\[\Large\rm 1=\frac{10B_c}{100+B_c}\]

zepdrix (zepdrix):

Multiply both sides by (100+Bc),\[\Large\rm 1(100+B_c)=10B_c\]\[\Large\rm 100+B_c=10B_c\] Understand how to solve it from there? Any steps confusing?

OpenStudy (anonymous):

I understand so far. next step is not sure.

zepdrix (zepdrix):

Combine like terms. Get the B's on the same side.

OpenStudy (anonymous):

Bc=10Bc-100?

zepdrix (zepdrix):

You moved the 100 to the other side.... You want your B's on the same side. You don't want to move the 100.

zepdrix (zepdrix):

Subtract B from each side.

OpenStudy (anonymous):

ok. 100-10Bc=Bc?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

???

zepdrix (zepdrix):

\[\Large\rm 100+B_c=10B_c\]Subtract B from each side,\[\Large\rm 100=10B_c-B_c\]

OpenStudy (anonymous):

100=9Bc!!

OpenStudy (anonymous):

ah. then Bc=9/100?

zepdrix (zepdrix):

yes!

OpenStudy (anonymous):

Thankk you soo much!!

zepdrix (zepdrix):

np c: